# Scaling Methods To Estimate Macroscopic Fundamental Diagrams in Urban Networks with Sparse Sensor Coverage

### v1

@@ Abstract
Accurately estimating traffic variables across unequipped portions of a network remains a significant challenge due to the limited coverage of sensor-equipped links, such as loop detectors and probe vehicles. A common approach is to apply uniform scaling, treating unequipped links as equivalent to equipped ones. This study introduces a novel framework to improve traffic variable estimation by integrating statistical scaling methods with geospatial imputation techniques. Two main approaches are proposed: (1) Statistical Scaling, which includes hierarchical and non-hierarchical network approaches, and (2) Geospatial Imputation, based on variogram modeling. The hierarchical scaling method categorizes the network into several levels according to spatial and functional characteristics, applying tailored scaling factors to each category. In contrast, the non-hierarchical method uses a uniform scaling factor across all links, ignoring network heterogeneity. The variogram-based geospatial imputation leverages spatial correlations to estimate traffic variables for unequipped links, capturing spatial dependencies in urban road networks. Validation results indicate that the hierarchical scaling approach provides the most accurate estimates, achieving reliable performance even with as low as 5% uniform detector coverage. Although the variogram-based method yields strong results, it is slightly less effective than the hierarchical scaling approach but outperforms the non-hierarchical method.
@@ 1 Introduction
The Macroscopic Fundamental Diagram (MFD) has emerged as a critical tool in network traffic management, providing a relationship between aggregate traffic variables such as flow, density, and speed across an entire urban road network [1]. This concept offers valuable insights for network traffic control [2, 3], particularly in perimeter control strategies [4, 5, 6], where traffic inflow and outflow are managed to optimize network-wide performance. By monitoring traffic at a macroscopic level, the MFD enables decision-makers to regulate traffic in real-time, improve congestion management, and enhance the overall efficiency of urban transportation systems.
@@
Despite its potential, estimating a reliable MFD for a city network requires extensive empirical data. Loop detector device (LDD), the most commonly used data source for MFD estimation, collects information from fixed sensors embedded in road infrastructure [7, 8, 9, 10, 4, 11, 12, 13, 14]. However, LDD faces limitations, such as positional biases in speed estimation, as well as the inability to install sensors on all network links, leading to incomplete coverage [13]. To address these limitations, floating car device (FCD), or Probe Vehicle Data, is often used in combination with LDD [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. FCD, derived from GPS-enabled vehicles, can capture vehicle trajectories and speeds, reducing the positional bias inherent in LDD. However, FCD has its own challenges, including the uncertain penetration rate of vehicles equipped with GPS in time and space. Estimating network-wide traffic flows based on the assumption of homogeneous FCD penetration introduces uncertainty in the flow estimation, particularly when coverage is sparse [28, 29, 30].
@@
The limitations of these data sources underscore a broader issue: the network-wide traffic data used for MFD estimation is typically incomplete. LDD and FCD are often available only for certain links in the network, resulting in what is known as an ‘equipped network MFD’, an MFD that only reflects the links where sensors or FCD are deployed [26]. Consequently, this equipped MFD may not be representative of the entire network. Estimating an accurate MFD for the entire city remains challenging due to the spatial coverage limitations of available data.
@@
To scale the equipped network MFD to the full city network, previous studies have attempted to partition the network into homogeneous areas and apply a scaling factor that adjusts for the ratio between the total length of links in the network and the length of the equipped links [26]. While this approach can offer some level of approximation, it suffers from significant limitations. The partitioning of the network based on equipped network data can lead to erroneous estimations of homogeneous areas, as the unequipped links may exhibit different traffic characteristics [31, 6, 23, 32, 33, 34]. Furthermore, applying a uniform scaling factor across areas with varying types of links, such as arterials and local roads, results in inaccurate flow estimates.
@@
Recent research has recognized the need for a more nuanced approach to spatial scaling of the MFD. To overcome the challenges associated with uniform scaling, a new methodology that accounts for the link hierarchy within a network is required. This method involves developing separate scaling factors for different hierarchical levels of the network, such as major arterials, collectors, and local streets. By accounting for the differences in traffic behavior at various link levels, this approach offers a more accurate estimation of network flow variables, allowing for the scaling of MFD from sensor-equipped portions of the network to the entire city.
This study proposes a comprehensive framework for spatial scaling of MFD that bridges the gap between sensor-equipped networks and the full urban road network. By incorporating link-hierarchical scaling and addressing the limitations of existing partitioning methods, we aim to improve the accuracy of network-wide MFD estimation, paving the way for more effective traffic management and control strategies at the city level.
@@ 1.1 Literature Review
1.1.1 Equipped Network
The literature about MFD estimation from ‘equipped network’ to ’full network’ can be broadly classified into two groups based on the data used: one using empirical data sources from the equipped network, the other relying on large-scale simulation frameworks for a city. The main challenges associated with the empirical data are the limited coverage of LDD and FCD throughout the network, and the simulation fails to replicate the network topology and link hierarchy of a complex network [28, 27]. Therefore, including simulation data can only ensure, to a specific extent, an understanding of the hierarchical links and their characteristics on estimation average network flow variables. Thus, we mainly focused on the existing empirical study on MFD estimation and calibration.
MFD estimation predominantly relied on LDD [15, 7] information on vehicle counts and occupancy rates on road segments, allowing the calculation of network-wide average flow and density. This approach was inspired by EDIE [35] definition of traffic flow metrics, where average flow and density are calculated based on the data collected from multiple equipped links within a network [28]. However, the reliance on LDD has certain limitations. For example, LDD is an Eulerian observation method, which may introduce bias in MFD estimation, particularly due to the spatial distribution of detectors [28]. The placement of detectors, often in congested areas or near traffic signals, can result in an overestimation of network density and contribute to scatter in the MFD [11]. However, a key challenge arises when MFD is estimated from LDD with only a subset of network links equipped with detectors. In such cases, the derived MFD reflects the equipped sub-network rather than the entire network [36, 26]. Despite this limitation, studies have shown that MFDs derived from a small fraction of links can still approximate critical density ranges observed in the full network MFD [11]. Mariotte et al. [26] introduced the concept of an ‘active network’ to improve MFD estimation, focusing on major arterials and secondary roads. They highlighted the importance of scaling the production MFD based on the ratio of total network length to equipped network length. The estimated scaling factor for production is based on the assumption of homogeneous network partitioning [37, 23, 32, 31, 34] using the equipped network information, leading to ambiguity in portioning for full network. Saffari et al. [38] expanded on this by investigating the selection of critical links for MFD estimation, though they noted that their findings were sensitive to the specific test bed used.
@@ 1.1.2 Spatial Scaling
The main challenge with imputing traffic states at the unequipped links is the estimation errors due lack of available information from the surrounding links. Network traffic state estimation approaches can be broadly classified into two main categories: model-driven and data-driven [39, 40]. High-accuracy model-driven methods often require additional information, such as detailed traffic signal settings or individual vehicle detector actuation data [41, 28]. While these methods have proven efficient in estimating traffic states, particularly at the link level [42, 43], their accuracy is highly dependent on the quality of the embedded models. Data-driven approaches leverage large amounts of historical data to train models using machine learning or statistical techniques [44, 45, 46, 40]. Since data-driven methods do not rely on strict theoretical assumptions, the accuracy of the models is largely influenced by the quantity and quality of the data. However, missing data is a common issue in data-driven methods, and several studies have attempted to address this problem [47, 48, 49]. Nowadays, graph convolution networks (GCN), generative adversarial networks (GANs), and physics-informed neural networks (PINNs) are gaining significant attention in estimating the missing traffic flow variables in unequipped networks. GCNs struggle with capturing dynamic temporal changes in traffic networks, as they are often designed for static graphs. Additionally, the hierarchical and complex topology of traffic networks complicates model training [50, 51, 52, 53, 54]. GANs face difficulties generating realistic traffic data due to the stochastic nature of traffic demand, making adversarial training unstable [55, 56]. PINNs, while incorporating physical laws, may struggle with noisy or incomplete data, as real-world traffic patterns can deviate significantly from theoretical models [57, 58]. Balancing accuracy and computational efficiency remains a significant challenge across these approaches.
1.1.3 Geo-spatial scaling
@@
In transportation data, kriging has been used to impute missing values in Annual Average Daily Traffic (AADT) datasets, with studies showing its superiority over traditional regression models like ordinary least squares (OLS) [59, 60]. The kriging method outperformed OLS, particularly in regions with moderate-to-high traffic volumes. However, under low traffic flow conditions, it tends to overestimate the missing data[60]. One of the limitations of the Kriging method is its reliance on Euclidean distance, which may not be ideal for transportation networks where distances are more appropriately measured along road networks. Zou et al. [61] proposed the use of an approximated road network distance based on isometric embedding theory to address this issue, improving the interpolation accuracy, particularly in complex road networks. Nevertheless, studies such as Selby and Kockelman [62] indicated that the use of road network distances does not always significantly improve prediction performance compared to traditional Euclidean distance-based methods.
@@
In addition, while kriging has demonstrated success in spatial imputation, its application to spatio-temporal data remains relatively underexplored. Some studies, such as those by Yang et al. [63], have extended kriging to account for both spatial and temporal dimensions. These efforts aim to better capture the spatio-temporal characteristics of traffic data. Moreover, Marcotte [64] proposed cokriging, which uses secondary correlated variables alongside the primary variable, has shown promise in various fields. The application of cokriging for traffic data imputation is an area worth exploring, especially given the availability of multiple sources of traffic data [65]. A recent study by Bae et al. [66] proposed the use of spatiotemporal cokriging to impute missing traffic flow speed data, integrating multiple data sources. Their results demonstrate that spatiotemporal cokriging improves the accuracy of imputation, particularly when missing data exhibit non-random patterns, such as blocks of missing data, which occur due to system malfunctions or maintenance issues. Later, Laval [67] showed that the traffic flow variables such as travel times, relaxation times, and delay in urban networks near critical density follow the scaling with total network length.
@@
Despite the advancements in using the MFD for traffic management, several gaps remain. Current methods rely on incomplete data from LDD and FCD, leading to inaccurate MFD estimates that don’t fully represent the entire network. Uniform scaling approaches overlook the hierarchical nature of urban road networks, resulting in imprecise flow and density estimates across different road types. Additionally, advanced data-driven models like GCNs and PINNs struggle with missing or noisy data, limiting their effectiveness in estimating traffic states. Moreover, traditional spatial scaling techniques such as kriging are not well-explored for complex road networks. A more refined approach, incorporating link-hierarchical scaling and better data imputation, is needed to improve MFD accuracy for complete network estimation.
2 Objectives and Contributions
Estimating a reliable MFD for an entire urban network remains a significant challenge due to the incomplete and biased coverage of traditional data sources like LDD and FCD. Existing methods often rely on uniform scaling approaches or oversimplified assumptions, which fail to capture the spatial and hierarchical variability of urban networks. These limitations hinder the accurate estimation of traffic flow dynamics across different road types, particularly in unequipped parts of the network, leading to suboptimal traffic management and control strategies. To address these challenges, this study develops a comprehensive framework for scaling network flow dynamics from equipped to full networks by incorporating road hierarchy and geospatial techniques.
@@ 2 Objectives and Contributions
    The primary contributions of this work include a novel hierarchical statistical scaling framework, which accounts for the distinct traffic characteristics of different road types, such as arterials and local streets, overcoming the inaccuracies of uniform scaling methods.
@@
    Adapts geospatial interpolation methods such as spatiotemporal variogram for traffic flow data, leveraging secondary data sources to address gaps and inconsistencies in LDD/FCD coverage and accurate estimation of network traffic dynamic.
@@
    Provides a robust evaluation of the proposed methodologies using real-world traffic data, demonstrating their effectiveness in capturing network-wide flow dynamics.
By addressing key challenges in spatial scaling, this study advances the state-of-the-art in MFD estimation, contributing to more efficient urban traffic management systems.
@@ 3 Methodology
The proposed methodology estimates traffic variables for an entire network based on partially sensor-equipped data using statistical scaling and geospatial imputation approaches. Initially, the network is classified into equipped and unequipped links. The statistical approach includes hierarchical and non-hierarchical scaling. In the hierarchical method, links are categorized into several levels based on spatial and functional characteristics (in this study: Hierarchy-1, Hierarchy-2, and Hierarchy-3), with tailored scaling factors applied to estimate traffic on unequipped links. The non-hierarchical method applies uniform scaling across all links. The geospatial approach employs variogram-based imputation to model spatial correlations between equipped and unequipped links, refining the traffic estimates. The output from both approaches generates an estimation of traffic variables across the whole network, incorporating data from both sensor-equipped links and estimated values for non-equipped links. Figure 1 represents an overview of the proposed methodologies in a flowchart.
@@
3.1 Statical Scaling

### v2

@@ Abstract
Accurately estimating traffic variables across unequipped portions of a network remains a significant challenge due to the limited amount of sensors-equipped links, such as loop detectors and probe vehicles. A common approach is to apply uniform scaling, treating unequipped links as equivalent to equipped ones, which leads to a strong bias in MFD estimation. Two main approaches are proposed: (1) Hierarchical Network Scaling and (2) Variogram-based data imputation. The hierarchical scaling method categorizes the network into several clusters according to spatial and functional characteristics, applying tailored scaling factors to each category. The variogram-based imputation leverages spatial correlations to estimate traffic variables for unequipped links, capturing spatial dependencies in urban road networks. Validation results show that the hierarchical scaling approach yields the most accurate estimates, demonstrating reliable performance with as little as 5% uniform detector coverage, while the variogram model fails to provide estimates. While the variogram-based method provides strong results with over 10% detector coverage, it is slightly less effective than the hierarchical scaling approach but performs better than the baseline non-hierarchical method.
@@ 1 Introduction
The Macroscopic Fundamental Diagrams (MFDs) has emerged as a critical tool in network traffic management, providing a relationship between aggregate traffic variables such as flow, density, and speed across an entire urban road network (Daganzo and Geroliminis, 2008). This concept offers valuable insights for network traffic control (Keyvan-Ekbatani et al., 2015; Ampountolas et al., 2017), particularly in perimeter control strategies (Aboudolas and Geroliminis, 2013; Mariotte and Leclercq, 2019; Jiang and Keyvan-Ekbatani, 2023), where traffic inflow and outflow are managed to optimize network-wide performances. By monitoring traffic at a macroscopic level, the MFDs enable decision-makers to regulate traffic in real time, improve congestion management, and enhance the overall efficiency of urban transportation systems.
@@
Despite its potential, estimating reliable MFDs for a city network requires extensive empirical data. Loop detector device (LDD), the most commonly used data source for MFDs estimation, collects information from fixed sensors embedded in road infrastructure (Buisson and Ladier, 2009; Geroliminis and Sun, 2011; Keyvan-Ekbatani et al., 2012; Saberi and Mahmassani, 2012; Aboudolas and Geroliminis, 2013; Keyvan-Ekbatani et al., 2013; Ambühl et al., 2021; Lee et al., 2023; Mousavizadeh and Keyvan-Ekbatani, 2024). However, LDD faces limitations, such as positional biases in speed estimation (Leclercq et al., 2014; Maiti, 2024), and uneven distribution of sensors in a network (Lee et al., 2023), leading to incomplete coverage. To address these limitations, floating-car devices (FCD), or probe is often used alongside LDD to enhance spatial coverage, primarily for improving spatial speed estimation (Geroliminis and Daganzo, 2008; Gayah and Daganzo, 2011; Geroliminis and Sun, 2011; Mahmassani et al., 2013; Tsubota et al., 2014; Ambühl and Menendez, 2016; Yang Beibei et al., 2016; Du et al., 2016; Saeedmanesh and Geroliminis, 2016; Ambühl et al., 2017; Dakic and Menendez, 2018; Mariotte et al., 2020a, b). However, FCD has its own challenges, including the unknown and variable penetration rate of vehicles equipped with GPS in time and space. Estimating network-wide traffic flows based on the assumption of homogeneous FCD penetration introduces uncertainty in the flow estimation, particularly when coverage is sparse (Leclercq et al., 2014; Shim et al., 2019; Fu et al., 2020).
@@
The limitations of these data sources underscore a broader issue: the network-wide traffic data used for MFD estimation is typically incomplete. LDD and FCD are often available only for certain links in the network, resulting in what is known as an ‘equipped network MFD’, an MFD that only reflects the links where sensors or FCD are deployed (Mariotte et al., 2020a). Consequently, this equipped MFD may not be representative of the entire network. This is critical for simulation studies as the demand is usually given for the full region, and estimating capacity and other variables on the equipped network only creates a mismatch with trips that could totally or partially happen in the non-equipped network.
@@
To scale the equipped network MFD to the full city network, previous studies have attempted to partition the network into homogeneous areas and apply a scaling factor that adjusts for the ratio between the total length of links in the network and the length of the equipped links, e.g., Mariotte et al. (2020a). While this approach can offer some level of approximation, it suffers from significant limitations. The partitioning of the network based on equipped network data can lead to erroneous estimations of homogeneous areas, as the unequipped links may exhibit different traffic characteristics (Jiang et al., 2023; Jiang and Keyvan-Ekbatani, 2023; Saeedmanesh and Geroliminis, 2016, 2017; Gu and Saberi, 2019; Johari et al., 2023). Furthermore, applying a uniform scaling factor across areas with varying types of links, such as arterials and local roads, may lead to inaccurate flow estimates.
@@
There is a need for a more advanced approach to spatial scaling of the MFD. To overcome the challenges associated with uniform scaling, a new methodology that accounts for the link hierarchy within a network is proposed. This method involves developing separate scaling factors for different hierarchical levels of the network, such as major arterials, collectors, and local streets. By accounting for the differences in traffic behavior at various link levels, this approach offers a more accurate estimation of network flow variables, allowing for accurate scaling of MFD from sensor-equipped portions of the network to the entire city.
@@ 1.1 Literature Review
1.1.1 Equipped Networks’ MFDs
The literature about MFD estimation from ‘equipped network’ to ‘full network’ can be broadly classified into two groups based on the data used: one using empirical data sources from the equipped network, the other relying on large-scale simulation frameworks for a city. The main challenges associated with the empirical data are the limited coverage of LDD and FCD throughout the network, while the simulations fail to replicate the multimodal interactions and traffic dynamics of a complex network (Leclercq et al., 2014; Mariotte et al., 2020b; Maiti and Chilukuri, 2023). Thus, we mainly focused on the existing empirical study on MFD estimation and calibration.
MFD estimation predominantly relied on LDD (Geroliminis and Daganzo, 2008; Buisson and Ladier, 2009) information on vehicle counts and occupancy rates on road segments, allowing the calculation of network-wide average flow and density. This approach was inspired by Edie (1963) definition of traffic flow metrics, where average flow and density are calculated based on the data collected from multiple equipped links within a network (Leclercq et al., 2014). However, the reliance on LDD has certain limitations. LDD is an Eulerian observation method, which may introduce bias in MFD estimation, particularly due to the spatial distribution of detectors (Leclercq et al., 2014). The placement of detectors, often in congested areas or near traffic signals, can result in an overestimation of network density (Lee et al., 2023; Maiti, 2024). Despite this limitation, studies have shown that MFDs derived from a small fraction of links can still approximate critical density ranges observed in the full network MFD (Keyvan-Ekbatani et al., 2013). Mariotte et al. (2020a) introduced the concept of an ‘active network’ to improve MFD estimation, focusing on major arterials and secondary roads. They highlighted the importance of scaling the production MFD based on the ratio of total network length to equipped network length. The estimated scaling factor for production is based on the assumption of homogeneous network partitioning (Zheng and Geroliminis, 2013; Saeedmanesh and Geroliminis, 2016, 2017; Jiang et al., 2023; Johari et al., 2023) using the equipped network information, leading to ambiguity in portioning for full network. Saffari et al. (2020) expanded on this by investigating the selection of critical links for MFD estimation, though they noted that their findings were sensitive to the specific test bed used.
There are two main options for estimating the traffic states in non-equipped networks to derive complete MFDs. The first involves grouping non-equipped links with clusters of links exhibiting similar behavior and determining a scaling factor for each cluster. The second approach focuses on imputing missing link variables for non-equipped links. The following review outlines the existing studies for spatial scaling and imputation in network traffic state estimation.
@@ 1.1.2 Spatial Scaling
The main challenge with imputing traffic states at the unequipped links is the estimation errors due to a lack of available information from the surrounding links and heterogeneity of traffic states among links. Network traffic state estimation approaches can be broadly classified into two main categories: model-driven and data-driven (Seo et al., 2017; Takayasu et al., 2022). High-accuracy model-driven methods often require additional information, such as detailed traffic signal settings or individual vehicle detector actuation data (Wu et al., 2011; Leclercq et al., 2014). While these methods have proven efficient in estimating traffic states, particularly at the link level (Nantes et al., 2016; Maiti and Chilukuri, 2024; Maiti et al., 2024; Maiti and Chilukuri, 2023), their accuracy is highly dependent on the quality of the embedded models. Data-driven approaches leverage large amounts of historical data to train models using machine learning or statistical techniques (Qing-Jie Kong et al., 2009; Anuar et al., 2015; Lu et al., 2018; Takayasu et al., 2022; Maiti and Chilukuri, 2023). Since data-driven methods do not rely on strict theoretical assumptions, the accuracy of the models is largely influenced by the quantity and quality of the data. However, missing data is a common issue in data-driven methods, and several studies have attempted to address this problem (Chen et al., 2003; Xu et al., 2015; Offor et al., 2019). Nowadays, graph convolution networks (GCN), generative adversarial networks (GANs), and physics-informed neural networks (PINNs) are gaining significant attention in estimating the missing traffic flow variables in unequipped networks. GCNs struggle with capturing dynamic temporal changes in traffic networks, as they are often designed for static graphs. Additionally, the hierarchical and complex topology of traffic networks complicates model training (Rossi et al., 2021; Taguchi et al., 2021; Castro-Correa et al., 2023; Bao et al., 2023; Zhang et al., 2023). GANs face difficulties generating realistic traffic data due to the stochastic nature of traffic demand, making adversarial training unstable (Zhang et al., 2019; Zheng et al., 2022). PINNs, while incorporating physical laws, may struggle with noisy or incomplete data, as real-world traffic patterns can deviate significantly from theoretical models (Zhang et al., 2024; Usama et al., 2022). Balancing accuracy and computational efficiency remains a significant challenge across these approaches.
1.1.3 Geo-spatial imputation
@@
In transportation data, kriging has been used to impute missing values in Annual Average Daily Traffic (AADT) datasets, with studies showing its superiority over traditional regression models like ordinary least squares (OLS) (Eom et al., 2006; Wang and Kockelman, 2009). The kriging method outperformed OLS, particularly in regions with moderate-to-high traffic volumes. However, under low traffic flow conditions, it tends to overestimate the missing data (Wang and Kockelman, 2009). One of the limitations of the Kriging method is its reliance on Euclidean distance, which may not be ideal for transportation networks where distances are more appropriately measured along road networks. Zou et al. (2012) proposed the use of an approximated road network distance based on isometric embedding theory to address this issue, improving the interpolation accuracy, particularly in complex road networks. Nevertheless, studies such as Selby and Kockelman (2013) indicated that the use of road network distances does not always significantly improve prediction performance compared to traditional Euclidean distance-based methods.
@@
In addition, while kriging has demonstrated success in spatial imputation, its application to spatio-temporal data remains relatively underexplored. Some studies, such as those by Yang et al. (2014), have extended kriging to account for both spatial and temporal dimensions. These efforts aim to better capture the spatio-temporal characteristics of traffic data. Moreover, Marcotte (1991) proposed cokriging, which uses secondary correlated variables alongside the primary variable, has shown promise in various fields. The application of cokriging for traffic data imputation is an area worth exploring, especially given the availability of multiple sources of traffic data (Shamo et al., 2015). A recent study by Bae et al. (2018) proposed the use of spatiotemporal cokriging to impute missing traffic flow speed data, integrating multiple data sources. Their results demonstrate that spatiotemporal cokriging improves the accuracy of imputation, particularly when missing data exhibit non-random patterns, such as blocks of missing data, which occur due to system malfunctions or maintenance issues. Later, Laval (2023) showed that the traffic flow variables such as travel times, relaxation times, and delay in urban networks near critical density follow the scaling with total network length.
@@
Despite the advancements in using the MFD for traffic management, several gaps remain. Current methods rely on incomplete data from LDD and FCD, leading to inaccurate MFD estimates that do not fully represent the entire network. Uniform scaling approaches overlook the hierarchical nature of urban road networks, resulting in imprecise flow and density estimates across different road types. Additionally, advanced data-driven models like GCNs and PINNs struggle with missing or noisy data, limiting their effectiveness in estimating traffic states. Moreover, traditional spatial scaling techniques such as kriging are not well-explored for complex road networks. A more refined approach, incorporating link-hierarchical scaling and better data imputation, is needed to improve MFD accuracy for complete network estimation. This study develops a comprehensive framework for scaling network flow dynamics from equipped to full networks by incorporating road hierarchy and spatial imputation techniques. The main contributions of this study are listed below:
@@
    This work includes a novel hierarchical network scaling framework, which accounts for the distinct traffic characteristics of different road types, such as arterials and local streets, overcoming the inaccuracies of uniform scaling methods.
@@
    Utilizes geospatial imputation methods, such as variogram models, for network traffic data imputation. It leverages secondary data sources to fill gaps and inconsistencies in LDD coverage, ensuring a more accurate estimation of network traffic dynamics.
@@
    Provides validation of the proposed methodologies using real-world traffic data, demonstrating their effectiveness in capturing network-wide traffic dynamics.
@@ 2 Methodology
The proposed methodology estimates traffic variables for an entire network based on partially sensor-equipped data using hierarchical network scaling and variogram-based imputation approaches. Initially, the network is classified into equipped and unequipped links. In the hierarchical approach, links are categorized into several hierarchical levels based on spatial and functional characteristics, with tailored scaling factors applied to estimate traffic on unequipped links on each hierarchy. The spatial imputation approach employs variogram-based imputation to model spatial correlations between equipped and unequipped links, refining the traffic estimates. The outputs from these methods provide comprehensive estimates of traffic variables across the network, integrating observed data with imputed values for unequipped links. The results are benchmarked against a non-hierarchical uniform scaling method, which applies consistent scaling factors across all links regardless of their spatial or functional characteristics. A summary of the methodological framework is provided in Figure 1.
@@
2.1 Network Variables Estimation and Scaling Issues
2.1.1 Uniform Scaling or Baseline